# Can Uncertainty Quantification Enable Better Learning-based Index Tuning?

### v1

@@ 4.4.1 Model Training
Compared to the original BE model, our BE-UQ model incorporates two additional modules: the decoder and MCD. Since the decoder cannot be trained concurrently with the original BE model, a redesigned training process is required. In our index advisor, the BE-UQ model is trained in two distinct phases. First, the BE model—which includes the feature extractor, hidden vector encoder, and Predictor—is trained following the original BE model’s method. In the second phase, the parameters of the feature extractor and hidden vector encoder are frozen, and the decoder is trained by minimizing the MSE loss.
@@
The rationale behind this training method is that BE remains the primary objective of the BE-UQ model, while UQ is secondary. Our approach ensures that the accuracy of the BE model is maintained, regardless of whether UQ is utilized. By preserving the accuracy of the hidden vector encoder, we ensure that the input information is effectively encoded in the intermediate representation $V^{h}$. This, in turn, enables the decoder to accurately reconstruct the input vector set $\hat{V}^{1}$ and correctly quantify uncertainty.
Although there are alternative ways to train the BE-UQ model, such as jointly training the hidden vector encoder and decoder followed by training the Predictor, this method shifts the encoder’s focus toward UQ rather than BE, which conflicts with our objectives. Therefore, our proposed training method is better suited for the index tuning task.
4.4.2 Uncertainty-Driven Model Updating
The performance of the BE-UQ model may decline when the database or query workload changes. To maintain the accuracy and stability of the model, it must be updated promptly when performance starts to decrease. There are two common ways to determine whether a machine learning model requires updating.
@@ 4.4.2 Uncertainty-Driven Model Updating
The first category samples predictions during model inference and compares them with actual values. A model update is needed if the error is significantly large. However, this approach has two main drawbacks. First, the actual benefits $B(q,\mathcal{I}_{0},\mathcal{I})$ for the sampled instances $(q,\mathcal{I}_{0},\mathcal{I})$ must be known. To determine $B(q,\mathcal{I}_{0},\mathcal{I})$, the indexes in $\mathcal{I}$ must be created, and the query $q$ must be executed based on those indexes, which incurs substantial time and space costs. Second, this method may cause delayed updates due to sampling errors.
The second category triggers model updates when a decrease in system efficiency is detected. However, this approach causes updates to lag behind the onset of performance regression. More importantly, in the context of index tuning, the optimal index is unknown. Even if an erroneous BE result leads to the creation of a suboptimal index, system performance may still improve slightly. Thus, it is difficult to determine whether the model requires updating based on this result.
@@
In the Beauty framework, we have a more accurate and timely metric to determine whether an update is needed: if the what-if tool estimates benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$, this likely indicates that the BE-UQ model no longer fits the current dataset, workload, or system conditions and requires updating. Additionally, the uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ returned by the model can help guide the update process.
@@
(1) If the what-if tool estimates the benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{1}>\theta_{1}$, this suggests that the BE-UQ model cannot provide effective latent vectors under the current workload. For example, if there have been substantial changes in query patterns or index configurations, the BE-UQ model may fail to produce accurate results. In this case, more data must be collected from the current workload to retrain or fine-tune the model.
@@
(2) If the what-if tool estimates the benefits for a significant portion of test instances $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{2}>\theta_{2}$, this indicates that the BE-UQ model is unable to distinguish between too many queries and index configurations in the current workload. Due to insufficient modeling information or flaws in feature extraction, the model estimates a wide range of benefits for the same test instance, leading to unreliable predictions. In such cases, the designer must identify the missing information or flaws in feature extraction and redesign the model based on the current workload.
4.4.3 Improving Efficiency.
@@
Therefore, when the feature extraction module is effective and the encoded information is sufficient, $\mathcal{U}_{2}$ quantification can be disabled. While this may slightly reduce the accuracy of the UQ, it significantly improves the inference efficiency of the BE-UQ model.
5 Empirical Study on Uncertainty Quantification
In this section, we compare different UQ methods from three aspects, namely the UQ capability, the efficiency and the impact on the accuracy of BE.
5.1 Datasets Construction
Three famous database benchmarks are adopted in the empirical study. Table II summarizes the databases and the workloads.
TPC-H: Following Leis et al. [38, 39], we randomly generated 30 queries from each template in TPC-H, excluding templates 2, 17, and 20. These were excluded because their queries have significantly longer execution times, which would disproportionately affect the overall workload cost. Indexes accelerating these queries are typically preferred over those for other templates.
TPC-DS: We also randomly generated 30 queries from each template in TPC-DS. Templates 4, 6, 9, 10, 11, 32, 35, 41, and 95 were excluded for the same reason given above.
JOB: This workload consists of 30 queries randomly generated from each template in JOB.
To increase the complexity of index tuning, we ingested synthetic queries into these workloads using the method devised by Yu et al. [39]. The enhanced workloads are referred to as TPC-H+, TPC-DS+ and JOB+, respectively.
@@ 5.1 Datasets Construction
Benchmark DB SF/Size # Queries # Templates # Synthetic Queries Avg. # Plan Nodes Maximum Runtime (mins)
TPC-H SF = 10 570 19 0 14 20
TPC-DS SF = 1 2700 90 0 29 45
JOB 13GB 3390 113 0 24 60
TPC-H+ SF = 10 950 19 380 10 30
TPC-DS+ SF = 1 4500 90 1800 19 60
JOB+ 13GB 5650 113 2260 16 75
To the best of our knowledge, no systematic experiments have previously explored uncertainty involved in index tuning tasks. To investigate the impacts of UQ on the accuracy and the efficiency of learning-based benefit estimators, as well as their ability to quantify uncertainty, we designed the following experimental methodology:
We ran the Autoadmin and Extend algorithms on three distinct workloads: TPC-H+, TPC-DS+, and JOB+, under various budget constraints to collect index configurations for tuning. After materializing these configurations, we gathered query execution data, represented as [(query, index config) $\rightarrow$ benefit], and formed a comprehensive dataset. This dataset was then divided into three segments based on query templates and indexes:
Training Dataset: Used for model training, containing approximately 50% of the total data. This subset only includes some of the query templates and indexes.
Test Dataset:Comprises about 15% of the total data, where each query or index configuration appeared in the training set but was not simultaneously seen by the model. This segment tests the model’s generalization capability within in-domain data.
Uncertainty Dataset:This dataset accounts for about 35% of the total data. It includes queries and indexes that are not present in the first two datasets, and no similar queries (from the same template) appear. For each model mentioned in Subsubsection 5.2.2, we will test using these samples, excluding those with errors below the 90th percentile of the Training Dataset errors. This way, the remaining samples are those the model fails to predict correctly due to U1-U4, and the model should assign a high uncertainty value to them.These samples are designated as Uncertainty Samples.
There are two points to note. First, the samples that different models fail to predict accurately vary, resulting in an inconsistent number of Uncertainty Samples, but the difference does not exceed 10%. Second, we use query costs to calculate the benefit rather than the execution time of the query, so we do not include U5-type uncertainty samples here.
5.2 Experimental Setup
5.2.1 BE Models
@@
In this experiment, we adopt two well-known learning-based BE models:
@@
(1) LIB. The BE model proposed in [7] encodes the operators and information in a query plan that can be influenced by indexes, forming a variable-length Index Optimizable Operations Set. After encoding through Multi-head Attention, the pooling operation retrieves $v^{h}$, culminating in the output of the BE via a MLP.
@@
(2) AMA. The model in [6] utilizes multiple channels to extract information and employs a weighted sum to compare query plans, selecting the faster option. Following [7], we replaced AMA’s classification layer with an MLP for benefit estimation.
It is important to note that while we aim to maintain consistency with the original paper regarding the two BE models, complete consistency may not be achievable. Our primary focus is on studying the impact of the UQ method rather than directly comparing the two BE models.
5.2.2 UQ Methods
@@
(1) Ours. This model integrates the hybrid quantification method proposed in this paper, resulting in LIB-Hybrid and AMA-Hybrid. We also compare an alternative version that utilizes efficiency enhancements, represented as LIB-AE and AMA-AE.
@@
(3) MCD. We compare three variants of this method in the experiment [40]: the first variant replaces all dropout layers in the model with MCD, resulting in LIB-MCD-all and AMA-MCD-all. The second variant applies MCD only to the last layer of the model’s Prediction Module, forming LIB-MCD-last and AMA-MCD-last. The third variant adopts the approach proposed in [37], which utilizes determinantal point processes for sampling to enhance diversity with minimal inference time increase, forming LIB-MCD-dpp and AMA-MCD-dpp.
@@
5.2.3 Implementation
@@
We set the relative benefit $rb=\frac{B(q,\mathcal{I}_{0},\mathcal{I})}{c(q,\mathcal{I}_{0})}$ as the prediction target. To maintain compatibility with the LIB framework, we exclude data entries that exhibit negative benefits from our dataset. For each model discussed, we generate thirty sets of hyperparameters and train on the training dataset. Subsequently, we retain the model that demonstrates the best performance based on evaluations using the test dataset.
@@
When quantifying uncertainty, we use the variance of results from five models as the quantification of uncertainty for the ensemble-based model. For other models, we perform inference 20 times [37] for each batch sample, using variance as the quantification of uncertainty.
(a) TPC-H+
(b) TPC-DS+
(c) JOB+
Figure 8: Uncertainty distribution of models using BNN, MCD, Ensemble, and HSTO.
Figure 9: Uncertainty distribution of models using Hybrid.
Figure 10: Uncertainty distribution of models using AE.
5.2.4 Hardware & Software
@@
The experiments were carried out on a Ubuntu server with two Intel Xeon 4210R CPUs (10 cores, 2.40GHz), 256GB of main memory, and one NVIDIA GeForce RTX 3060 GPU. The DBMS is PostgreSQL 12.13.
5.3 Results on Uncertainty Quantification
We present Raincloud plots illustrating the uncertainty distributions for sixteen models across various benchmarks. Figure 8 displays the plots for models based on BNN, MCD, Ensemble, and HSTO. Figures 9 and 10 show the plots for models based on Hybrid and AE, respectively. Each subplot depicts the uncertainty distribution of samples as calculated by the model for the training dataset, test dataset, and Uncertainty Samples. Consequently, the greater the uncertainty of the orange samples representing Uncertainty Samples in the graph and the higher their differentiation from other samples, the more precisely the model’s uncertainty modeling is demonstrated, enabling it to effectively distinguish uncertainty data. The red lines in the graphs represent the average uncertainties across the three datasets.
TABLE III: Percentage of Uncertainty Samples exceeding the uncertainty threshold set by the Training Dataset.
LIB-based Model TPC-H+ TPC-DS+ JOB+ AMA-based Model TPC-H+ TPC-DS+ JOB+
BNN 13.03 40.20 16.04 BNN 45.26 65.32 53.67
MCD-all 68.08 40.98 26.22 MCD-all 51.73 54.17 45.61
MCD-last 61.79 29.32 5.92 MCD-last 21.76 21.27 22.11
MCD-dpp 16.02 36.91 8.35 MCD-dpp 78.28 34.83 30.76
Ensemble 79.01 53.44 29.32 Ensemble 38.73 77.62 60.22
AE 71.85 31.13 30.93 AE 92.68 84.19 61.00
Hybrid 81.80 57.65 32.88 Hybrid 91.74 78.13 47.82
$\mathcal{U}_{1}$ 51.88 27.21 20.68 $\mathcal{U}_{1}$ 88.87 54.21 20.00
$\mathcal{U}_{2}$ 59.70 43.46 20.76 $\mathcal{U}_{2}$ 54.93 47.50 34.87
STO 68.41 34.86 23.51
HSTO 63.64 35.13 24.58
Table III further illustrates the filtering capabilities of various methods for Uncertainty Samples. Since Hybrid includes two UQ modules, we adopt the following approach to ensure comparability with other models: for $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, we select the 90th percentile of the uncertainty distribution from the Training Dataset as their corresponding thresholds. We calculate the proportion $P$ of samples in the Training Dataset with uncertainty values lower than these thresholds. If $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are probabilistically independent, this proportion should be 81%. For other models with only one UQ module, we employ the $P$-th percentile of the uncertainty distribution from the Training Dataset as a threshold. The table shows the proportion of samples in the Uncertainty Samples that exceed this threshold. The higher the proportion, the stronger the model’s ability to identify uncertain samples. For each benchmark in the table, we highlight the model with the most effective filtering in green and the second most effective in blue. Note that $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, as parts of the Hybrid, are excluded from this ranking.
From these results, we can draw the following conclusions:
First, the AE and Hybrid models consistently perform the best, followed by the Ensemble method. The other methods lag significantly behind in both stability and optimal performance in UQ. The best performance is seen with the AMA-AE model under the TPC-H+, which can identify 92.68% of Uncertainty Samples while misclassifying about 19% of training samples, significantly outperforming general-purpose methods. In more than two-thirds of the benchmarks, the AE and Hybrid models consistently rank among the top two, demonstrating better stability than the Ensemble method.

### v2

@@ IV-D1 Model Training
Compared to the original BE model, our BE-UQ model incorporates two additional modules: the decoder and the MCD. Since the decoder cannot be trained concurrently with the original BE model, the training process is redesigned. In particular, the BE-UQ model is trained in two distinct phases. In phase 1, the BE model—which includes the feature extractor, hidden vector encoder, and Predictor—is trained using the original method proposed for the BE model. In phase 2, the parameters of the feature extractor and the hidden vector encoder are frozen, and the decoder is trained to minimize the MSE loss. The framework and hyperparameters for the decoder can be selected independently of those for the encoder.
@@
The rationale behind this training method is that BE is the primary objective of the BE-UQ model, while UQ is secondary. Our training method ensures the accuracy of BE, regardless of whether UQ is enabled. By preserving the accuracy of the hidden vector encoder, we guarantee that the input information is effectively encoded in the intermediate representation $V^{h}$. This, in turn, enables the decoder to accurately reconstruct the input vector set $\hat{V}^{1}$ and correctly quantify uncertainties.
IV-D2 Uncertainty-Driven Model Updating
The performance of the BE model may degrade when the underlying database or query workload undergoes changes. To maintain the model’s accuracy and stability, it is imperative to promptly initiate updates when a decline in performance is detected. There are two primary approaches to determining when a model requires an update.
@@ IV-D2 Uncertainty-Driven Model Updating
The first approach involves sampling predictions during model inference and comparing them to actual values. An update is triggered if the error exceeds a predefined threshold. However, this method necessitates the creation of indexes in $\mathcal{I}$ to obtain the actual benefits $B(q,\mathcal{I}_{0},\mathcal{I})$, thereby incurring substantial time and space costs. Additionally, this approach may result in delayed updates due to sampling errors.
The second approach triggers model updates upon detection of a decrease in system efficiency. However, this method suffers from a significant drawback: it causes updates to lag behind the onset of performance regression.
@@
In the Beauty framework, we have a more accurate and timely metric to determine whether an update is needed: if the what-if tools estimate benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$, this likely indicates that the BE-UQ model no longer fits the current dataset, workload, or system conditions and requires updating. Additionally, the uncertainties $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ returned by the model can help guide the update process.
@@
(1) If the what-if tools estimate the benefits for a significant portion of test samples $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{1}>\theta_{1}$, this suggests that the BE-UQ model cannot provide effective latent vectors under the current workload. For example, if there have been substantial changes in query patterns or index configurations, the BE-UQ model may fail to produce accurate results. In this case, more data must be collected from the current workload to retrain or fine-tune the model.
@@
(2) If the what-if tools estimate the benefits for a significant portion of test instances $(q,\mathcal{I}_{0},\mathcal{I})$ due to $\mathcal{U}_{2}>\theta_{2}$, this indicates that the BE-UQ model is unable to distinguish between too many queries and index configurations in the current workload. Due to insufficient modeling information or flaws in feature extraction, the model estimates a wide range of benefits for the same test instance, leading to unreliable predictions. In such cases, the designer must identify the missing information or flaws in feature extraction and redesign the model based on the current workload.
IV-D3 Efficient Version.
@@
Therefore, in scenarios where high efficiency is required, the quantification of $\mathcal{U}_{2}$ can be disabled. As will be shown in the evaluation in section V, eliminating the quantification of $\mathcal{U}_{2}$ can significantly increase the efficiency of model inference, while slightly decreasing the accuracy of UQ.
V Empirical Study on Uncertainty Quantification
In this section, we integrate UQ methods into the BE model, thereby establishing BE-UQ models. We will systematically evaluate each BE-UQ model across three dimensions: uncertainty quantification capability, benefit estimation accuracy, and inference efficiency.
@@ V Empirical Study on Uncertainty Quantification
Benchmark DB SF/Size # Queries # Templates
# Synthetic
Queries
Avg. #
Plan Nodes
Maximum
Runtime
TPC-H+ SF = 10 950 19 380 10 90
TPC-DS+ SF = 10 4500 90 1800 19 120
JOB+ 13GB 5650 113 2260 16 180
V-A Construction of Datasets
We prepared databases and query workloads based on three famous benchmarks, TPC-H, TPC-DS, and JOB. For each benchmark, we create a workload that consists of two groups of queries. (1) For each query template provided in the benchmark, we randomly generated $30$ queries. Following the procedure commonly conducted in previous studies [32, 17, 33], we exclude templates 2, 17, and 20 from TPC-H and exclude templates 4, 6, 9, 10, 11, 32, 35, 41, and 95 from TPC-DS. These templates were excluded because queries generated using these templates have much longer execution times than queries generated using other templates. If these templates were used, index tuning tends to recommend indexes that only accelerate queries generated using these templates. (2) To increase the difficulty of benefit estimation and index tuning, complex synthetic queries that include joins or nested queries were randomly generated. To ensure each synthetic query to yield at least one result, we adopted the method [33] in which the results are generated first, followed by the parameters in the query. We call the extended benchmarks with synthetic queries TPC-H+, TPC-DS+, and JOB+, respectively. Their statistics are given in Table II.
To the best of our knowledge, there have been no comprehensive experimental studies on the performance of UQ methods integrated with learning-based BE models in the context of index tuning. Therefore, we design the following experiments.
First, for each benchmark, we construct a comprehensive dataset $D$ that consists of records of the form (query, indexes, benefit). In particular, we use the algorithms AutoAdmin [11] and Extend [19] to generate candidate index configurations for the query workload. For each generated index configuration $\mathcal{I}$, we run each query $q$ in the workload before and after creating the indexes in $\mathcal{I}$ and record the actual benefit $b$ that these indexes bring to $q$, thus obtaining a record $(q,\mathcal{I},b)$ in $D$.
To evaluate the performance of UQ methods integrated with learning-based BE models, we divide $D$ into $3$ disjoint partitions $D_{train}$, $D_{test}$ and $D_{eval}$, where $D_{train}$ is used to train the BE-UQ model, $D_{test}$ is used to test the BE accuracy of the BE-UQ model, and $D_{eval}$ is used to evaluate the UQ capability of the BE-UQ model. Notably, we deliberately select $D_{eval}$ to exhibit substantially distinct distributional characteristics compared to $D_{train}$ and $D_{test}$. Due to the distinction in the distributions of $D_{train}$ and $D_{eval}$, the BE-UQ model trained on $D_{train}$ is supposed to give high uncertainties for the samples in $D_{eval}$. Therefore, we can use the fraction of samples in $D_{eval}$ that are identified as unreliable by the BE-UQ model using the method proposed in subsubsection IV-C4 as a metric of UQ performance.
In particular, $D_{eval}$ is constructed as follows. First, randomly select some query templates and columns from the benchmark. Then, pick the records $(q,\mathcal{I},b)$ in $D$ into $D_{eval}$ if $q$ is generated from a selected template, or $\mathcal{I}$ contains the selected columns. We carefully select templates and columns so that $D_{train}$, $D_{test}$ and $D_{eval}$ approximately contain $50\%$, $15\%$, and $35\%$ of records in $D$, respectively.
V-B Experimental Setup
V-B1 BE Models
@@
In this experiment, we adopt two well-known learning-based BE models.
@@
(1) LIB. The LIB model [7] encodes the operators and information in a query plan that can be influenced by indexes, forming a variable-length Index Optimizable Operations Set. After encoding through the Multi-head Attention mechanism, the pooling operation retrieves $v^{h}$, culminating in the output of BE via an MLP.
@@
(2) AMA. The AMA model [6] utilizes multiple channels to extract information from the plans provided by what-if tools and employs a weighted sum to compare these plans, ultimately selecting the faster option. Following [7], we replaced AMA’s classification layer with an MLP for BE.
V-B2 UQ Methods
@@
(1) Ours. This model integrates the hybrid quantification method proposed in this paper, resulting in LIB-Hybrid and AMA-Hybrid. Furthermore, we compare the efficiency-enhanced variants outlined in  subsubsection IV-D3, denoted as LIB-AE and AMA-AE.
@@
(3) MCD. We compare three variants of this method in the experiment [34]: the first variant replaces all dropout layers in the model with MCD, resulting in LIB-MCD-all and AMA-MCD-all. The second variant applies MCD only to the last layer of the model’s Prediction Module, forming LIB-MCD-last and AMA-MCD-last. The third variant adopts the approach proposed in [31], which utilizes determinantal point processes for sampling to enhance diversity, forming LIB-MCD-dpp and AMA-MCD-dpp.
@@
V-B3 Implementation
We define the relative benefit as the prediction target:
$\frac{B(q,\mathcal{I}_{0},\mathcal{I})}{c(q,\mathcal{I}_{0})}.$ (4)
@@ V-B3 Implementation
To maintain compatibility with the LIB framework, we exclude data entries that exhibit negative benefits from the training dataset. For each model discussed, we generate thirty sets of hyperparameters and train on the training dataset. Subsequently, we retain the model that demonstrates the best performance based on evaluations using the test dataset.
Initially, the BE-UQ models constructed using our method implemented a mirror-symmetric decoder architecture with respect to their corresponding encoders, adhering to established conventions in this domain. However, when this symmetric configuration failed to achieve satisfactory convergence in the decoder component, we preserved the fundamental mechanisms (e.g., attention) while restructuring the neural architecture. Specifically, we systematically varied both the neuron density per layer and the network depth, thereby generating five distinct decoder variants. Following comprehensive retraining of each decoder configuration, we retained exclusively the variant that exhibited superior performance metrics.
@@
When quantifying uncertainty, we use the variance of results from five models as the quantification of uncertainty for the ensemble-based model. For other models, we perform inference $20$ times [31] for each batch sample, using variance as the quantification of uncertainty. The uncertainty threshold is established following the method described in subsubsection IV-C4, with the parameter $\alpha$ set to $1.3$.
Figure 8: Uncertainty distributions produced by various BE-UQ models for the TPC-H+ benchmark.
V-B4 Hardware & Software
@@
The experiments were carried out on a Ubuntu server with two Intel Xeon 4210R CPUs (10 cores, 2.40GHz), 256GB of main memory and one NVIDIA GeForce RTX 3060 GPU. The DBMS is PostgreSQL 12.13.
V-C Performance of Uncertainty Quantification
TABLE III: Proportion of unreliable samples in $D_{eval}$ successfully identified by various UQ methods.
LIB-based
Model
TPC-H+ TPC-DS+ JOB+
AMA-based
Model
TPC-H+ TPC-DS+ JOB+
BNN 12.08 5.41 1.41 BNN 3.39 71.13 38.17
MCD-all 67.95 39.16 20.60 MCD-all 19.30 69.33 22.36
MCD-last 60.95 11.79 1.67 MCD-last 10.51 14.73 7.31
MCD-dpp 7.74 5.54 2.16 MCD-dpp 76.50 11.25 13.20
Ensemble 75.30 60.77 24.86 Ensemble 7.72 50.94 53.10
AE 76.21 62.71 24.33 AE 77.01 76.90 53.75
Hybrid 83.69 74.38 30.68 Hybrid 91.40 89.37 59.81
STO 68.81 48.81 18.21
HSTO 62.50 39.33 19.08
In this subsection, we evaluate the uncertainty quantification (UQ) performance of various BE-UQ models.
Evaluation Metric. We infer all BE-UQ models on $D_{train}$, $D_{test}$, and $D_{eval}$, obtaining the BE error and uncertainty of each sample in these datasets. We establish the BE error threshold for determining the reliability of a model’s BE results as the 90th percentile of BE errors on $D_{train}$. Samples in $D_{eval}$ with BE errors exceeding this threshold are designated as unreliable samples.
To quantitatively evaluate the UQ performance of a BE-UQ model, we use the proportion of unreliable samples that are identified as unreliable estimates, that is, with uncertainties exceeding the specified uncertainty threshold, as the evaluation metric. A higher proportion indicates a better recall of unreliable samples realized by the BE-UQ model, thereby a higher UQ capability of the model.
Experimental Results. The experimental results are presented in Table III. For each benchmark, the most effective BE-UQ model is highlighted in green, and the second most effective one is highlighted in blue. Furthermore, Figure 8 shows the uncertainty distributions produced by various BE-UQ models on $D_{train}$, $D_{test}$, and $D_{eval}$ using Raincloud plots. In each Raincloud plot, the horizontal axis is the uncertainty value, and the red line connects the means of uncertainty distributions. Due to space constraints, we only illustrate the results for the TPC-H+ benchmark; the results for other benchmarks are available in our open-source repository. From these results, we derive the following conclusions:
(1) The Hybrid and AE models consistently outperform other methods, followed by the Ensemble approach. The remaining methods exhibit significantly inferior performance in both stability and unreliable sample detection capabilities. The Hybrid method achieves superior results across all benchmarks, particularly within the TPC-H+ benchmark, where the AMA-Hybrid model successfully identifies more than $91\%$ of unreliable samples, substantially surpassing alternative approaches. In more than five-sixths of evaluation scenarios, the AE framework demonstrates performance exceeding that of the Ensemble method, exhibiting superior stability. HSTO, despite being specifically engineered for Self-Attention mechanisms, fails to demonstrate enhanced effectiveness in this particular task. Among the three MCD-based methodologies, MCD-all demonstrates superior efficacy, achieving optimal results in five of the six evaluation scenarios.
(2) Figure 8 shows that the uncertainties of the samples in $D_{train}$ mostly follow long-tailed distributions, which is consistent with the claim in subsubsection IV-C4 that the uncertainties follow long-tailed distributions. This verifies the rationale behind our method for setting uncertain thresholds.